# A variational formulation of stochastic thermodynamics. Part I: Finite-dimensional systems

### v1

@@ I Introduction
For continuous-time systems, which are the focus of this paper, the Langevin equation must be complemented by an appropriate fluctuation-dissipation relation (FDR) to ensure thermodynamic consistency. Such an FDR will be shown to provide a well-defined interpretation of entropy production by connecting the first and second laws, as well as guaranteeing the existence of an equilibrium steady state for any closed system. However, there is no systematic approach for determining such relations in a general setting, and the Einstein relation [2] is often assumed by default [1], justified by the requirement that the corresponding Fokker–Planck equation is satisfied by the Boltzmann distribution in the steady state [3, 4]. Moreover, thermodynamically consistent modeling of complex systems remains challenging when accounting for stochastic dynamics, particularly in nonequilibrium regimes, such as active matter systems [5, 6]. We emphasize the meaning of FDRs as a generalization of the Einstein relation, i.e. the relation between dissipative forces and noise strength, and is not to be identified with fluctuation-dissipation or fluctuation-response theorems as in how the system responds to an external perturbation.
The goal of this paper is to present a novel approach to the thermodynamically consistent modeling of stochastic finite-dimensional, continuous-time systems. To this end, we extend the variational formulation of nonequilibrium thermodynamics presented in Ref. [7] to the stochastic regime.
@@
By extending this framework to the stochastic regime, we generalize it to describe systems driven far from equilibrium. The proposed variational principle for stochastic thermodynamics is constrained by the first and second laws of thermodynamics, under the assumption of time-scale separation. No steady-state condition is imposed, nor is the thermal bath required to be infinite. Instead, results from statistical mechanics are recovered in the appropriate limits, while the formulation reproduces macroscopic nonequilibrium thermodynamics in the vanishing-noise limit. Moreover, as shown in subsequent sections, this approach naturally extends key features of stochastic thermodynamics, including an individual trajectory level description of both configurational and thermal variables, and fluctuation theorems (FTs) in the thermodynamic phase space. Remarkably, once thermodynamic consistency is enforced via the FDRs, the equations for entropy production derived from the variational approach assume the same form as in standard ST, up to a reformulation in an extended phase space.
The paper is organized as follows. Section II provides an overview of the framework of ST, which can be skipped by readers already familiar with the topic, except perhaps for subsection II.4. Section III introduces the variational principle for ST and illustrates it with examples of (finite-dimensional) closed and open systems. Section IV offers a discussion of our findings, and Section V summarizes the main conclusions of this study. Additional analytical derivations are organized into five appendices at the end.
@@ II.1 Representations of Langevin processes
$\partial_{t}\mathcal{P}$ $=$ $-\partial_{x^{\nu}}\left(\frac{p^{\nu}}{m}\mathcal{P}\right)-\partial_{p^{\nu}}\left[\left(F^{\nu}-g^{\nu}_{k}\partial_{p^{\mu}}g^{\mu}_{k}\right)\mathcal{P}-g^{\nu}_{k}g^{\mu}_{k}\partial_{p^{\mu}}\mathcal{P}\right]$ (3)
$=$ $-\nabla_{\mathbf{x}}\cdot\mathbf{j}_{x}-\nabla_{\mathbf{p}}\cdot\mathbf{j}_{p},$
where $\mathcal{P}(\mathbf{x},\mathbf{p},t)$ denotes the probability density function (PDF) of a certain configuration $(\mathbf{x},\mathbf{p})$ at time $t$, and $\mathbf{j}$ the probability current. While the Langevin equation (1a)-(1b) provides dynamical information of an individual trajectory, the FP equation (3) provides the ensemble level information.
@@
For convenience, we denote the full phase-space variable as $\mathbf{z}=(\mathbf{x},\mathbf{p})$. The initial and final times are denoted by $t_{0}$ and $t$, respectively, with corresponding configurations $\mathbf{z}_{0}$ and $\mathbf{z}_{t}$. The path-probability can be defined as
@@
$K(\mathbf{z}_{t},t;\mathbf{z}_{0},t_{0})=\int_{\mathbf{z}_{0}}^{\mathbf{z}_{t}}\prod_{n=1}^{N-1}\prod_{\nu}\frac{dz_{n}^{\nu}\delta(m\dot{x}^{\nu}-p^{\nu})}{\det(g^{\nu\mu}_{n+1/2})\sqrt{4\pi\epsilon}}\,e^{-\mathcal{A}[\mathbf{z}]},$ (6)
@@
$\mathcal{D}\mathbf{z}=\mathcal{D}\mathbf{x}\mathcal{D}\mathbf{p}\delta^{(d)}(\mathbf{p}-m\mathbf{\dot{x}}).$ (7)
@@ II.2 Stochastic entropy production
In stochastic thermodynamics (ST), entropy production (EP) is defined as the breaking of time-reversal symmetry (TRSB), and is quantitatively measured by the Kullback-Leibler (KL) divergence—or statistical distance—between the probability distributions of a forward process and its time-reversed counterpart [19, 20, 21]:
$\langle\Delta s_{tot}\rangle\coloneq D_{KL}[P[\mathbf{z}]||P[\hat{\mathbf{z}}]]=\int_{\mathbf{z}_{0}}^{\mathbf{z}_{t}}\mathcal{D}\mathbf{z}P[\mathbf{z}]\text{ln}\frac{P[\mathbf{z}]}{P[\hat{\mathbf{z}}]}$, (9)
@@
The KL divergence is an invariant (under non-singular transformations), nonnegative flat operator [22], such that the definition (9) already incorporates the second law of ST $\langle\Delta s_{tot}\rangle\geq 0$. Moreover, as a consequence of this definition, the integral fluctuation theorem (IFT) follows [23]:
$\langle e^{-\Delta s_{tot}}\rangle=\int\mathcal{D}\mathbf{z}P[\mathbf{z}]\frac{P[\hat{\mathbf{z}}]}{P[\mathbf{z}]}=\int\mathcal{D}\mathbf{z}P[\hat{\mathbf{z}}]=1$ (10)
@@
$\Delta s_{tot}=\text{ln}\frac{P[\mathbf{z}]}{P[\hat{\mathbf{z}}]}=\Delta s_{m}+\Delta s,$ (11a)
$\Delta s_{m}\coloneq\text{ln}\frac{P[\mathbf{z}|\mathbf{z}_{0}]}{P[\hat{\mathbf{z}}|\mathbf{z}_{t}]},\quad\Delta s\coloneq\text{ln}\frac{\mathcal{P}(\mathbf{z}_{0},t_{0})}{\mathcal{P}(\mathbf{z}_{t},t)}.$ (11b)
@@
$\mathcal{S}(t)$ $\coloneq\langle s(t)\rangle=-\int_{\mathbb{R}^{2d}}d\mathbf{z}\,\mathcal{P}(\mathbf{z},t)\ln\mathcal{P}(\mathbf{z},t).$ (12b)
@@
$\Delta s_{m}$ $=\text{ln}\frac{P[\mathbf{z}|\mathbf{z}_{0}]}{P[\hat{\mathbf{z}}|\mathbf{z}_{t}]}=\text{ln}\frac{\mathcal{N}\{\mathbf{z}_{n}\}\,\exp\{-\mathcal{A}\}}{\mathcal{N}\{\mathbf{\hat{z}}_{n}\}\,\exp\{-\hat{\mathcal{A}}\}}=\hat{\mathcal{A}}-\mathcal{A}$ (17)
@@
Here, the backward action is given by $\hat{\mathcal{A}}[\mathbf{z},\lambda]=\mathcal{A}[(\mathbf{x}(u),-\mathbf{p}(u),\lambda(u))]$, with $u=t-\tau$. Since the integration limits coincide under this change of variables, we can relabel the dummy index $du=d\tau$, allowing direct subtraction of the integrands [24]. Taking the time derivative gives the medium EPR (see [25, §8.2.1] for a discussion of time integrals of Markovian variables):
@@
$\dot{s}_{tot}$ $=\dot{s}+\dot{s}_{m}=-\frac{\partial_{t}\mathcal{P}}{\mathcal{P}}-\frac{\partial_{x_{\mu}}\mathcal{P}}{\mathcal{P}}\frac{p^{\mu}}{m}+\frac{D_{\mu\nu}j^{\nu}_{d}}{\mathcal{P}}(\dot{p}^{\mu}-F^{\mu}_{+})$ (19)
@@
where $\mathbf{j}_{p}$ is the probability current along the momentum dimension defined in (3). See Appendix A for a proof of this relation. In the following derivations, we take into account the natural boundary conditions associated with the PDF [25]:
@@
$\dot{S}_{tot}\coloneq\langle\!\langle\dot{s}_{tot}\rangle\!\rangle=\int d\mathbf{z}\,\frac{j^{\mu}_{d}}{\mathcal{P}}D_{\mu\nu}j^{\nu}_{d},$ (23)
$\dot{S}_{m}\coloneq\langle\!\langle\dot{s}_{m}\rangle\!\rangle=\int d\mathbf{z}\,j^{\mu}_{d}D_{\mu\nu}(F^{\nu}_{-}-g^{\nu}_{k}\partial_{p_{\alpha}}g^{\alpha}_{k})-\langle\partial_{p_{\mu}}F^{\mu}_{+}\rangle,$ (24)
again consistent with [26] for additive noise. Notably, only the third term in Eq. (19) survives the averaging process. Provided the covariance matrix $D^{\nu\mu}$ is positive definite, the average total EPR yields a quadratic form and hence verifies $\dot{S}_{tot}\geq 0$, with equality being restricted to equilibrium where dissipative currents vanish $({\bf j}_{d}=\underline{0})$.
Moreover, $\dot{\mathcal{S}}=\langle\!\langle\dot{s}\rangle\!\rangle$, which can be shown by directly differentiating Shannon entropy (see [27, §A.3] for total time derivatives of integrals):
@@
$=\langle\!\langle\dot{s}_{tot}\rangle\!\rangle-\langle\!\langle\dot{s}_{m}\rangle\!\rangle=\langle\!\langle\dot{s}\rangle\!\rangle.$
@@ II.3 Stochastic energetics
Stochastic energetics deals with the stochastic extension of the first law and its connections with the second law obtained via ST. We illustrate it with the system Eq. (1b), but we distinguish three different kind of forces, $\mathbf{F}=-\nabla_{\mathbf{x}}V(\mathbf{x},\lambda)-\underline{\underline{\bm{\gamma}}}(\mathbf{x},\mathbf{p})\mathbf{p}/m+\mathbf{f}_{+}(\mathbf{x},\mathbf{p})$, where $V(\mathbf{x},\lambda)$ gives the potential energy, $\underline{\underline{\bm{\gamma}}}$ the damping coefficient and $\mathbf{f}_{+}$ non-conservative even forces, as could be an external manipulation by an agent [26] or feedback cooling [28]. By analogy with the previous section, we identify the reversible and irreversible contributions as $\mathbf{F}_{+}=-\nabla_{\mathbf{x}}V(\mathbf{x},\lambda)+\mathbf{f}_{+}$ and $\mathbf{F}_{-}=-\underline{\underline{\bm{\gamma}}}\cdot\mathbf{p}/m$. Following Sekimoto’s approach [8], we can derive the first law from the total energy of the system $E$:
@@
$=(-\underline{\underline{\bm{\gamma}}}(\mathbf{x},\mathbf{p})\cdot\mathbf{p}/m+\underline{\underline{\mathbf{g}}}(\mathbf{x},\mathbf{p})\circ\bm{\zeta})\cdot\frac{\mathbf{p}}{m}dt$
@@
$\dot{q}\coloneq(-\underline{\underline{\bm{\gamma}}}(\mathbf{x},\mathbf{p})\cdot\mathbf{p}/m+\underline{\underline{\mathbf{g}}}(\mathbf{x},\mathbf{p})\circ\bm{\zeta})\cdot\frac{\mathbf{p}}{m},$ (30a)
$\dot{w}\coloneq\mathbf{f}_{+}\cdot\frac{\mathbf{p}}{m}+\frac{\partial V}{\partial\lambda}\dot{\lambda}.$ (30b)
@@
In the particular scenario where $\underline{\underline{\bm{\gamma}}}(\mathbf{x},\mathbf{p})=\underline{\underline{\bm{\gamma}}}$ is a constant, the Einstein relation $T\underline{\underline{\bm{\gamma}}}=\underline{\underline{\mathbf{D}}}$, where $\underline{\underline{\mathbf{D}}}$ is the covariance matrix and $T$ is the absolute temperature, implies for Eq. (18):
$\dot{s}_{m}$ $=$ $-\frac{1}{T}(-\underline{\underline{\bm{\gamma}}}\cdot\mathbf{p}/m+\underline{\underline{\mathbf{g}}}\circ\bm{\zeta})\cdot\frac{\mathbf{p}}{m}-\nabla_{\mathbf{p}}\cdot\mathbf{f}_{+}$ (31)

### v2

# A variational formulation of stochastic thermodynamics: Finite-dimensional systems

@@ I Introduction
However, the extensions of the first and second laws to the stochastic regime in ST are not systematically connected. On one hand, entropy acquires a new definition from the perspective of information theory as a raw measure of irreversibility [73, 87, 78, 79, 64, 35], while stochastic energetics is approached from a different point of view, namely Sekimoto’s approach [80]. To reconcile the stochastic first and second laws, we require our model to satisfy thermodynamic consistency, providing a well-defined physical interpretation of entropy beyond its information-theoretic meaning. The principle of local detailed balance (LDB) offers precisely this constructive link, which fundamentally stems from the time-reversal symmetry of the microscopic laws of Physics or micro-reversibility, see Ref. [62].
For continuous-time systems, which are the focus of this paper, the principle of LDB manifests in the Langevin equation through the fluctuation-dissipation relation (FDR), in order to ensure thermodynamic consistency. We emphasize that throughout this work we refer to FDR as generalized fluctuation-dissipation Einstein relation linking noise and dissipation [33, 88, 74, 87], also termed FDR of the second kind, and is not to be identified with the FDR of the first kind as in how the system responds to an external perturbation [61]. Note that imposing LDB in the set-up of Markov dynamics implies an FDR of the second kind, which in turn implies a standard FDR of the first kind near equilibrium, see Ref. [61].
Thus, the principle of LDB is essential for constructing physically meaningful and thermodynamically interpretable models [64, 5, 62]. While there are different methods to derive FDRs, including fixing the equilibrium distribution in the Fokker-Planck steady state [95, 52] or imposing the LDB condition directly [62, 64], determining such relations in a general setting in the context of Langevin dynamics remains challenging, and the Einstein relation [30] is often assumed by default [79]. As a result, the model at hand often lacks a clear thermodynamic interpretation [62]. The goal of this paper is to present a novel general and systematic approach to the thermodynamically consistent modeling of stochastic finite-dimensional, Langevin systems. To this end, we extend the variational formulation of nonequilibrium thermodynamics presented in Ref. [44] to the stochastic regime, thereby establishing a principled connection between ST and analytical mechanics, in line with the growing interest [7].
@@
By extending this framework to the stochastic regime, we generalize it to describe systems driven far from equilibrium. The proposed variational principle for ST is built from the first and second laws of thermodynamics, under the assumptions of time-scale separation, large bath size and micro-reversibility. The FDR can be derived systematically in general settings without assuming a fixed steady state, without relying on problematic path probabilities, and without requiring thermal baths to be infinite. Instead, results from statistical mechanics are recovered in the appropriate limits, while the formulation reproduces macroscopic nonequilibrium thermodynamics in the vanishing-noise limit. Moreover, as shown in subsequent sections, this approach naturally extends key features of ST, including an individual trajectory level description of both configurational and thermal variables, and fluctuation theorems (FTs) in the thermodynamic phase space. Remarkably, once thermodynamic consistency is enforced via the FDRs, the equations for entropy production derived from the variational approach assume the same form as in standard ST, up to a reformulation in an extended phase space.
Here two remarks are in order. Firstly, in Ref. [61] it is illustrated how not in all physical cases LDB needs to be true, providing examples such as non-equilibrium baths or large system-bath/bath-bath couplings. The role of the variational approach in this regard will be assessed in Sec. IV. Secondly, as the variational principle is framed in Lagrangian mechanics, it only covers Langevin-type dynamics and not the more general setting of Markov chains.
The paper is organized as follows. Section II provides an overview of the framework of ST, which can be skipped by readers already familiar with the topic, except perhaps for subsection II.4. Section III introduces the variational principle for ST and illustrates it with examples of (finite-dimensional) closed and open systems. Section IV offers a discussion of our findings, and Section V summarizes the main conclusions of this study. Additional analytical derivations are organized into six appendices at the end.
@@ II.1 Representations of Langevin processes
$\partial_{t}\mathcal{P}$ $=-\partial_{x^{\nu}}\left(\frac{p^{\nu}}{m}\mathcal{P}\right)-\partial_{p^{\nu}}\left[\left(F^{\nu}-g^{\nu}_{k}\partial_{p^{\mu}}g^{\mu}_{k}\right)\mathcal{P}-g^{\nu}_{k}g^{\mu}_{k}\partial_{p^{\mu}}\mathcal{P}\right]$
$=-\nabla_{\mathbf{x}}\cdot\mathbf{j}_{x}-\nabla_{\mathbf{p}}\cdot\mathbf{j}_{p},$ (3)
where $\mathcal{P}(\mathbf{x},\mathbf{p},t)$ denotes the probability density function (PDF) of a certain configuration $(\mathbf{x},\mathbf{p})$ at time $t$, and $\mathbf{j}$ the probability current. While the Langevin equation (1a)-(1b) provides dynamical information of an individual trajectory, the FP equation (II.1) provides the ensemble level information.
@@
For convenience, we denote the full phase-space variable as $\mathbf{z}=(\mathbf{x},\mathbf{p})$. The initial and final times are denoted by $t_{0}$ and $t$, respectively, with corresponding configurations $\mathbf{z}_{0}$ and $\mathbf{z}_{t}$. The path probability can be defined as
@@
$K(\mathbf{z}_{t},t;\mathbf{z}_{0},t_{0})=\int_{\mathbf{z}_{0}}^{\mathbf{z}_{t}}\mathcal{D}\mathbf{z}\prod_{n=0}^{N-1}\frac{1}{\det(g^{\nu\mu}_{n+1/2})(4\pi\epsilon)^{d/2}}\,e^{-\mathcal{A}[\mathbf{z}]},$ (6)
@@
$\mathcal{D}\mathbf{z}=\mathcal{D}\mathbf{x}\mathcal{D}\mathbf{p}\delta^{(d)}(\mathbf{p}-m\mathbf{\dot{x}})=\lim_{\epsilon\rightarrow 0}\prod_{n=1}^{N-1}\prod_{\nu}dz_{n}^{\nu}\delta(m\dot{x}^{\nu}-p^{\nu}).$ (7)
@@ II.2 Stochastic entropy production
In ST, entropy production (EP) is defined as the breaking of time-reversal symmetry (TRSB), and is quantitatively measured by the Kullback-Leibler (KL) divergence—or statistical distance—between the probability distributions of a forward process and its time-reversed counterpart [71, 73, 87]:
$\langle\Delta s_{\rm tot}\rangle:=D_{KL}[P[\mathbf{z}]||P[\hat{\mathbf{z}}]]=\int_{\mathbf{z}_{0}}^{\mathbf{z}_{t}}\mathcal{D}\mathbf{z}P[\mathbf{z}]\text{ln}\frac{P[\mathbf{z}]}{P[\hat{\mathbf{z}}]}$, (9)
@@
The KL divergence is an invariant (under non-singular transformations), nonnegative flat operator [54], such that the definition (9) already incorporates the second law of ST $\langle\Delta s_{\rm tot}\rangle\geq 0$. Moreover, as a consequence of this definition, the integral fluctuation theorem (IFT) follows [78]:
$\langle e^{-\Delta s_{\rm tot}}\rangle=\int\mathcal{D}\mathbf{z}P[\mathbf{z}]\frac{P[\hat{\mathbf{z}}]}{P[\mathbf{z}]}=\int\mathcal{D}\mathbf{z}P[\hat{\mathbf{z}}]=1$ (10)
@@
$\Delta s_{\rm tot}=\ln\frac{P[\mathbf{z}]}{P[\hat{\mathbf{z}}]}=\Delta s_{m}+\Delta s,$ (11a)
$\Delta s_{m}:=\ln\frac{P[\mathbf{z}|\mathbf{z}_{0}]}{P[\hat{\mathbf{z}}|\mathbf{z}_{t}]},\quad\Delta s:=\ln\frac{\mathcal{P}(\mathbf{z}_{0},t_{0})}{\mathcal{P}(\mathbf{z}_{t},t)}.$ (11b)
@@
$\mathcal{S}(t)$ $:=\langle s(t)\rangle=-\int_{\mathbb{R}^{2d}}d\mathbf{z}\,\mathcal{P}(\mathbf{z},t)\ln\mathcal{P}(\mathbf{z},t).$ (12b)
@@
$\Delta s_{m}$ $=\ln\frac{P[\mathbf{z}|\mathbf{z}_{0}]}{P[\hat{\mathbf{z}}|\mathbf{z}_{t}]}=\text{ln}\frac{\mathcal{N}\{\mathbf{z}_{n}\}\,\exp\{-\mathcal{A}\}}{\mathcal{N}\{\mathbf{\hat{z}}_{n}\}\,\exp\{-\hat{\mathcal{A}}\}}=\hat{\mathcal{A}}-\mathcal{A}$ (17)
@@
Here, the backward action is given by $\hat{\mathcal{A}}[\mathbf{z},\lambda]=\mathcal{A}[(\mathbf{x}(u),-\mathbf{p}(u),\lambda(u))]$, with $u=t-\tau$. Since the integration limits coincide under this change of variables, we can relabel the dummy index $du=d\tau$, allowing direct subtraction of the integrands [89]. Taking the time derivative gives the medium EPR (see [75, §8.2.1] for a discussion of time integrals of Markovian variables):
@@
$\dot{s}_{\rm tot}$ $=\dot{s}+\dot{s}_{m}=-\frac{\partial_{t}\mathcal{P}}{\mathcal{P}}-\frac{\partial_{x_{\mu}}\mathcal{P}}{\mathcal{P}}\frac{p^{\mu}}{m}+\frac{D_{\mu\nu}j^{\nu}_{d}}{\mathcal{P}}(\dot{p}^{\mu}-F^{\mu}_{+})$ (19)
@@
where $\mathbf{j}_{p}$ is the probability current along the momentum dimension defined in (II.1). See Appendix A for a proof of this relation. In the following derivations, we take into account the natural boundary conditions associated with the PDF [75]:
@@
$\dot{S}_{\rm tot}:=\langle\!\langle\dot{s}_{\rm tot}\rangle\!\rangle=\int d\mathbf{z}\,\frac{j^{\mu}_{d}}{\mathcal{P}}D_{\mu\nu}j^{\nu}_{d},$ (23)
$\dot{S}_{m}:=\langle\!\langle\dot{s}_{m}\rangle\!\rangle=\int d\mathbf{z}\,j^{\mu}_{d}D_{\mu\nu}(F^{\nu}_{-}-g^{\nu}_{k}\partial_{p_{\alpha}}g^{\alpha}_{k})-\langle\partial_{p_{\mu}}F^{\mu}_{+}\rangle,$ (24)
again consistent with [55] for additive noise. Notably, only the third term in Eq. (19) survives the averaging process. Provided the covariance matrix $D^{\nu\mu}$ is positive-definite (PD), the average total EPR yields a quadratic form and hence verifies $\dot{S}_{\rm tot}\geq 0$, with equality being restricted to equilibrium where dissipative currents vanish $({\bf j}_{d}=\bm{0})$.
Moreover, $\dot{\mathcal{S}}=\langle\!\langle\dot{s}\rangle\!\rangle$, which can be shown by directly differentiating Shannon entropy (see [84, §A.3] for total time derivatives of integrals):
@@
$=\langle\!\langle\dot{s}_{\rm tot}\rangle\!\rangle-\langle\!\langle\dot{s}_{m}\rangle\!\rangle=\langle\!\langle\dot{s}\rangle\!\rangle.$
@@ II.3 Stochastic energetics
Stochastic energetics deals with the stochastic extension of the first law and its connections with the second law obtained via ST. We illustrate it with the system Eq. (1b), but we distinguish three different kind of forces, $\mathbf{F}=-\nabla_{\mathbf{x}}V(\mathbf{x},\lambda)-\underline{\underline{\boldsymbol{\gamma}}}(\mathbf{x},\mathbf{p})\mathbf{p}/m+\mathbf{f}_{+}(\mathbf{x},\mathbf{p})$, where $V(\mathbf{x},\lambda)$ gives the potential energy, $\underline{\underline{\boldsymbol{\gamma}}}$ the damping coefficient and $\mathbf{f}_{+}$ non-conservative even forces, as could be an external manipulation by an agent [55] or feedback cooling [68]. By analogy with the previous section, we identify the reversible and irreversible contributions as $\mathbf{F}_{+}=-\nabla_{\mathbf{x}}V(\mathbf{x},\lambda)+\mathbf{f}_{+}$ and $\mathbf{F}_{-}=-\underline{\underline{\boldsymbol{\gamma}}}\cdot\mathbf{p}/m$. Following Sekimoto’s approach [80], we can derive the first law from the total energy of the system $E$:
@@
$=(-\underline{\underline{\boldsymbol{\gamma}}}(\mathbf{x},\mathbf{p})\cdot\mathbf{p}/m+\underline{\underline{\mathbf{g}}}(\mathbf{x},\mathbf{p})\circ\boldsymbol{\zeta})\cdot\frac{\mathbf{p}}{m}dt$
@@
$\dot{q}:=(-\underline{\underline{\boldsymbol{\gamma}}}(\mathbf{x},\mathbf{p})\cdot\mathbf{p}/m+\underline{\underline{\mathbf{g}}}(\mathbf{x},\mathbf{p})\circ\boldsymbol{\zeta})\cdot\frac{\mathbf{p}}{m},$ (30a)
$\dot{w}:=\mathbf{f}_{+}\cdot\frac{\mathbf{p}}{m}+\frac{\partial V}{\partial\lambda}\dot{\lambda}.$ (30b)
@@
In the particular scenario where $\underline{\underline{\boldsymbol{\gamma}}}(\mathbf{x},\mathbf{p})=\underline{\underline{\boldsymbol{\gamma}}}$ is a constant, the Einstein relation $T\underline{\underline{\boldsymbol{\gamma}}}=\underline{\underline{\mathbf{D}}}$, where $\underline{\underline{\mathbf{D}}}$ is the covariance of the random forces and $T$ is the absolute temperature, implies for Eq. (18):
$\dot{s}_{m}$ $=$ $-\frac{1}{T}(-\underline{\underline{\boldsymbol{\gamma}}}\cdot\mathbf{p}/m+\underline{\underline{\mathbf{g}}}\circ\boldsymbol{\zeta})\cdot\frac{\mathbf{p}}{m}-\nabla_{\mathbf{p}}\cdot\mathbf{f}_{+}$ (31)